how does the phase decrease as the wave moves in the direction.delta of phase angle is directly proportional to delta of displacement
phase angle=-k x where k is propagation constant and -undicates decrease in phase in the direction of wave.plz explain

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The phase of a wave refers to a sinusoidal function such as the following:

x(t) = A\cdot \cos( 2 \pi f t + \theta )\,
y(t) = A\cdot \sin( 2 \pi f t + \theta ) = A\cdot \cos( 2 \pi f t + \theta -\pi/2),\,

where A, f, and \scriptstyle \theta are constant parameters. These functions are periodic with period \scriptstyle T = 1/f, and they are identical except for a displacement of \scriptstyle T/4 along the \scriptstyle t axis. The term phase can refer to several different things:

  • It can refer to a specified reference, such as \scriptstyle \cos( 2 \pi f t)\,, in which case we would say the phase of \scriptstyle x(t) is \scriptstyle \theta, and the phase of \scriptstyle y(t) is \scriptstyle \theta -\pi/2.
  • It can refer to \scriptstyle \theta, in which case we would say \scriptstyle x(t) and \scriptstyle y(t) have the same phase but are relative to different references.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi